OpenStudy (anonymous):

Solve for x:(1/a+b+x)=(1/a)(1/b)+(1/x) with steps

OpenStudy (anonymous):

Like this?\[\frac{1}{a}+b+x=\frac{1}{a} \times \frac{1}{b}+\frac{1}{x}\]

OpenStudy (anonymous):

Or this...\[\frac{1}{a+b+x} = \frac{1}{a \times b} + \frac{1}{x}\]?

OpenStudy (anonymous):

b=1/a and x=-2/a+2

OpenStudy (anonymous):

I am assuming it's mine one... \[abx = (a+b+x)(ab+x)\] \[ x^2+(a+b)x+(a+b)ab=0\] Now you can solve it.